OpenStudy (anonymous):

write 10/20 in lowest terms

undeadknight26 (undeadknight26):

ok so: 10 divided my 20 equals .5 and .5 converted to fraction is?

undeadknight26 (undeadknight26):

its 1/2.

OpenStudy (texaschic101):

divide both number by 10 and what do you get ?

OpenStudy (anonymous):

im sorry that wasnt the question its 2/6 in lowest terms

OpenStudy (danielbarriosr1):

1/3

OpenStudy (anonymous):

thankz

OpenStudy (hesan):

10 /20 = 1/2 = 0.5

OpenStudy (texaschic101):

you need to find a number that will go into both numbers equally. Then divide each number by that number

OpenStudy (danielbarriosr1):

Np

OpenStudy (hesan):

if you divide 10 and 20 by the same number i.e 2 then you will get 1/2 i.e equal to 0.5

OpenStudy (hesan):

It will never ever become 1/3

OpenStudy (danielbarriosr1):

it is 2/6 in lowest terms @Hesan So it is 1/3

OpenStudy (hesan):

it becomes 1/2 not 2/6

undeadknight26 (undeadknight26):

@Hesan 10/20 was the wrong question...2/6 was the right question.

OpenStudy (danielbarriosr1):

@quecanchero wrote: im sorry that wasnt the question its 2/6 in lowest terms

OpenStudy (danielbarriosr1):

Thanks @undeadknight26

OpenStudy (hesan):

Ooo I see....

OpenStudy (hesan):

Then you are wright...